# Computing Wasserstein Barycenters through Gradient Flows

### v2

@@ C.3 Domain Adaptation
Appendix D Missing Proofs and Derivations
D.1 Connecting equations 12 and 16
In this section, we the empirical flow in equation 16 from the continuity equation. As a reminder, the continuity equation is,
$\partial_{t}P_{t}=\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})).$
Here we want to plug $\hat{P}_{t}=n^{-1}\sum_{i=1}^{n}\delta_{z_{t,i}}$ into this equation. On the one hand, plugging the formula for the empirical measure into the r.h.s. of this equation gives us,
$\text{div}(\hat{P}_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t}))=\text{div}\biggr(\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\nabla(\delta\mathbb{F}(\hat{P}_{t}))(z_{t,i})\biggr).$ (45)
Here, we used the definition for the Wasserstein gradient (c.f., equation 36), i.e., $\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla(\delta\mathbb{F}(P))(z_{t,i})$. On the other hand, due to the linearity of derivatives and the chain rule,
$\partial_{t}\hat{P}_{t}=\dfrac{1}{n}\sum_{i=1}^{n}\partial_{t}\delta_{z_{t,i}^{(P)}}=-\dfrac{1}{n}\sum_{i=1}^{n}\text{div}(\partial_{t}z_{t,i}\delta_{z_{t,i}})=\text{div}\biggr(-\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\partial_{t}z_{t,i}\biggr).$ (46)
Equating the r.h.s. of equations 45 and 46 gives us the continuous-time update $\partial_{t}z_{t,i}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t})$. We can discretize this differential equation in time, using the Forward Euler scheme,
$z_{\tau+1,i}=z_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}),$
where $\alpha>0$ is the time-step and $\tau$ is the discrete time index.
D.2 Proposition 3.1
As a recap, in this section we provide a formal proof that when,
$P=\sum_{i}\pi_{i}^{(P)}(\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}}),$
and under the ground cost,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime}),$
the $\mathbb{MW}_{2}$ distance becomes,
$\mathbb{MW}_{2}(P,Q)^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\omega^{\star}_{ij}(\mathbb{W}_{2}(P_{i},Q_{j})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}).$
As a reminder, we have the following proposition,
Proposition D.1.
Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be 2 GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$. Let the ground cost $c$ be,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
Proof.
Our first step is decomposing the Wasserstein distance $\mathbb{W}_{2}(P_{i},Q_{j})^{2}$ into $\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}$. Indeed,
$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int c(z,z^{\prime})d\gamma(z,z^{\prime}),$
where $z=(x,y)$ is the feature-label pair. Note that since $P_{i}=P_{i,x}\otimes\delta_{\nu_{i}^{(P)}}$, $\gamma$ is supported on the set $\{(x,y,x,y^{\prime}):y=\nu_{i}^{(P)},y^{\prime}=\nu_{j}^{(Q)}\}$. Thus, the OT plan has the structure:
$\gamma=\gamma_{x}\otimes\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}$
where $\gamma_{x}$ is the OT plan between Gaussians $P_{i,x}$ and $Q_{j,x}$. This means that the previous integral can be decomposed into,
$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int(\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2})d\gamma(z,z^{\prime}),$
which we can decompose two terms. The first,
$\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma(z,z^{\prime})$ $=\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma_{x}(x,x^{\prime})=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2},$
and,
$\int\int\rho(y,y^{\prime})^{2}d\gamma(z,z^{\prime})$ $=\int\int\rho(y,y^{\prime})^{2}\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}dydy^{\prime}=\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2},$
which gives the desired decomposition. ∎
The result in the previous proposition stems from the special structure we defined for the GMMs. Since the result is agnostic with respect the label metric, it holds for $\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})=\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}$
D.3 Theorem 3.1
  1. 1.
    The exponential convergence given by the PL inequality,
    $\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (47)
  2. 2.
    An uniform bound on $P\mapsto\mathbb{B}_{\mathcal{Q}}(P)$ controlling the error of approximating $Q_{k}$ with $\hat{Q}_{k}$.
These results are available, respectively, in [chewi2024statistical] and [portales2025sample]. For completeness, we re-state and demonstrate them. We start with a modified version of [chewi2024statistical, Corollary 5.17],
Proposition D.2.
(Measure theoretic Gronwall’s Lemma [chewi2024statistical]) Let $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ be a functional that satisfies the PL inequality (c.f. 47) with constant $C_{\text{PL}}>0$, then,
Proof.
Define $\phi(t)=\mathbb{F}(P_{t})-\mathbb{F}^{\star}$. Then,
$\partial_{t}\phi(t)$ $=\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$
From Grönwall’s Lemma [chewi2024statistical, Lemma 5.16],
We now present the uniform approximation bound of [portales2025sample] for empirical barycenters,
Theorem D.1.
(Barycenter Empirical Approximation error [portales2025sample]) Let $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ be a finite family of measures on $\mathcal{P}_{p,\text{ac}}(\mathcal{B}(0,R))$ for $R>0$. Let $\{z_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}$, $i=1,\cdots,n_{k}$ be a set of $K$ i.i.d. samples from each $Q_{k}$. Let $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}$. Let $p\in[1,+\infty)$. Then,
$\mathbb{E}\biggr[\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{sup}}\,|\mathbb{B}_{\mathcal{Q}}(\hat{P})-\widehat{\mathbb{B}}_{\mathcal{Q}}(\hat{P})|\biggr]\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$ (49)
where,
[…]

### v3

@@ E.3 Wasserstein Barycenter Transport
E.4 Adapting Neural Network Barycenter Solvers to DA
Here, we cover how we adapted previous neural network-based barycenters to perform domain adaptation. There is one main challenge with using unlabeled barycenters for domain adaptation: we cannot directly move the barycenter to the target domain, as we would end up with more unlabeled data. A straightforward solution to this problem is moving the available labeled data that we have, that is, the source domain data. For this scenario, the barycenter domain works as a pivot domain, as we transport the source to the target through the composition,
$T_{Q_{k}\rightarrow Q_{T}}(x_{i}^{(Q_{k})})=T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})).$ (67)
In comparison, by computing a labeled barycenter, we synthesize data in $P^{\star}$.
As we demonstrated in Tables 3 through 7, this is generally beneficial for domain adaptation. In this way, we can acquire a labeled dataset in the target domain through $\{\{T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})),y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$. To isolate the effect of the quality of the computed barycenters, we estimate the mappings as we did in equations 65 and 66, that is,
$T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})=n\sum_{j=1}^{n}\gamma_{k,i,j}^{\star}x_{j}^{(P)}\quad T_{P\rightarrow Q_{k}}(x_{i}^{(P)})=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$ (68)
Appendix F Additional Experiments
In total, we compare five domain adaptation benchmarks, divided into computer vision, neuroscience and chemical engineering. An overview of the benchmarks is presented in Table 1 in our main paper. Here, we give further details on each of the benchmarks. For computer vision, we use the Office 31 saenko2010adapting and Office Home venkateswara2017deep benchmarks. For neuroscience, we use the BCI-CIV-2a brunner2008bci and ISRUC khalighi2013automatic benchmarks. For chemical engineering, we use the TEP benchmark montesuma2024benchmarking.
F.1 Evaluation Protocol and Backbones
As we describe in the following, all benchmarks deal with multi-class classification. We measure performance using the standard classification accuracy,
For the Office 31, Office Home, Tennessee Eastman Process, and BCI-CIV-2a, we adopt the so-called leave-one-domain-out evaluation. For example, given $3$ domains $\{A,B,C\}$, we perform $3$ experiments, leaving one domain as the target domain at each time. Then, one of these experiments would be $\{A,B\}\rightarrow C$, where $\{A,B\}$ are the sources, and $C$ is the target.
As we discuss in the main paper, we perform adaptation over embeddings. To that end, we use the same experimental setting as montesuma2024lighter for the Office 31, Office Home, and TEP benchmarks. More specifically, we train a ResNet 50 and ResNet 101 for the Office 31 and Office Home benchmarks, and a CNN for the TEP benchmark. For this latter backbone, we refer readers to montesuma2024benchmarking for further details. For neuroscience benchmarks, we follow the experimental protocol of wang2024cbramod, where we fine-tune CBraMod, a transformer-based foundation model, on available source domain data. We use the exact same training settings as these authors.
F.2 Fine-Grained Results
Algorithm $\mathcal{X}\times\mathcal{Y}$ A D W Avg. $\uparrow$
ResNet50 - 67.50 95.00 96.83 86.40
WJDOT turrisi2022multi - 67.77 97.32 95.32 86.80
DaDiL-E montesuma2023multi - 70.55 100.00 98.83 89.79
DaDiL-R montesuma2023multi - 70.90 100.00 98.83 89.91
GMM-DaDiL montesuma2024lighter - 72.47 100.00 99.41 90.63
Discrete cuturi2014fast 70.55 96.43 96.49 87.82
NormFlow visentin2025computing 65.67 97.32 94.73 85.91
CW2B korotin2021continuous 69.16 94.64 95.32 86.37
NOT kolesov2024estimating 69.33 92.85 96.49 86.22
U-NOT gazdieva2024robust 68.29 97.32 95.32 86.97
WGF (ours) 69.16 98.21 96.49 87.96
Discrete montesuma2023multi 67.94 98.21 97.66 87.93
GMM montesuma2024lighter 70.13 99.11 96.49 88.54
WGF (ours) 70.03 99.11 99.41 89.52
Table 3: Classification accuracy per domain for the Office 31 benchmark.
Office 31 saenko2010adapting contains, in total, 4652 images among 31 classes of objects. These images are divided into 3 different domains: Amazon, dSLR, and Webcam. Images in the Amazon domain represent products, usually with an homogeneous background, medium resolution, and studio lightning. There are 2817 images in this domain. The dSLR domain contains high-resolution images obtained through a digital SLR camera in an office environment. There are 498 images in this domain. Finally, the Webcam domain contains images of objects captured with a webcam, also in an office environment. There 795 images in this domain. As analyzed in ringwald2021adaptiope, this benchmark is one of the most commonly used in the literature. Despite its moderate size, it presents challenges of its own. For instance, since the Amazon domain images were collected automatically, it has label noise.
We show our results in the Office 31 benchmark in Table 3. Overall, dictionary learning methods such as DaDiL montesuma2023multi or GMM-DaDiL montesuma2024lighter remain state-of-the-art for this benchmark. However, our gradient flow algorithm is able to improve previous methods, closing the gap between these two kinds of algorithms. In comparison with unsupervised barycenters, using labels in the ground-cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) has a clear advantage.
Algorithm $\mathcal{X}\times\mathcal{Y}$ Ar Cl Pr Rw Avg. $\uparrow$
ResNet101 - 72.90 62.20 83.70 85.00 75.95
WJDOT turrisi2022multi - 74.28 63.80 83.78 84.52 76.59
DaDiL-E montesuma2023multi - 77.16 64.95 85.47 84.97 78.14
DaDiL-R montesuma2023multi - 75.92 64.83 85.36 85.32 77.86
GMM-DaDiL montesuma2024lighter - 77.16 66.21 86.15 85.32 78.81
Discrete cuturi2014fast 76.13 63.91 83.11 85.55 77.17
NormFlow visentin2025computing 77.16 62.54 83.10 84.63 76.86
CW2B korotin2021continuous 74.69 63.46 82.54 85.09 76.44
NOT kolesov2024estimating 72.22 63.57 82.77 82.91 75.36
U-NOT gazdieva2024robust 75.30 64.94 82.88 84.28 76.85
WGF (ours) 74.89 64.60 83.89 85.21 77.56
Discrete montesuma2023multi 76.54 63.91 83.78 84.74 77.25
GMM montesuma2024lighter 75.31 64.26 86.71 85.21 77.87
WGF (ours) 76.13 65.52 85.81 86.24 78.42
Table 4: Classification accuracy per domain for the Office-Home benchmark.
Office-Home venkateswara2017deep contains, 15,500 images among 65 classes of objects. There are four domains: Art, Clipart, Product, and Real-World. The Art domain contains artistic depictions of objects, such as sketches, paintings, and ornamentation. It has 2427 images. The Clipart domain is a collection of graphic art representing the objects. This domain has 4365 images. Product contains images of objects without a background. It has 4439 images. Real-World contains images of objects captured with a camera. It contains 4357 images. In this sense, the Product domain is similar to the Amazon domain in the Office 31 benchmark, and the Real-World is similar to dSLR or Webcam. In comparison with the Office31 benchmark, the Office-Home benchmark adds another factor of domain variation: style (e.g., Art vs. Real-World). We show our results in Table 4. In comparison with the Office 31 benchmark, our methods get even closer to the state-of-the-art. Again, using labels in the ground cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) offers a clear advantage.
Remark F.1.
We follow previous works on visual domain adaptation, especially saenko2010adapting and venkateswara2017deep, and report the classification accuracy on fixed partitions of the target domains. For that reason, Tables 3 and 4 do not have confidence intervals. While it would be possible to run these experiments with random partitions (e.g., $k$-fold cross validation), this would mean that our results would not be comparable to previous established results in montesuma2024lighter, for instance.
Tennessee-Eastman Process (TEP) [reinartz2021extended, montesuma2024benchmarking] is a benchmark in chemical engineering. This benchmark comprises a set of simulations of a chemical reaction that occurs inside a reactor. In total, 34 variables are measured over time. As a result, we have a set of time-series representing how the reaction progresses over each simulation. In some of these simulations, one of 28 possible faults can be introduced. The goal is, based on the readings, to determine which kind of fault, or its absence, has occurred. As a result, there are 29 classes. Each simulation is also associated with a production mode, which drastically changes the statistical properties of the data generation process, thus introducing a distribution shift. As a result, each domain corresponds to one of these modes of operation. We refer readers to montesuma2024benchmarking and reinartz2021extended for more details.
CNN - 80.82 $\pm$ 0.96 63.69 $\pm$ 1.71 87.47 $\pm$ 0.99 79.96 $\pm$ 1.07 74.44 $\pm$ 1.52 84.53 $\pm$ 1.12 78.48
WJDOT turrisi2022multi - 89.06 $\pm$ 1.34 75.60 $\pm$ 1.84 89.99 $\pm$ 0.86 89.38 $\pm$ 0.77 85.32 $\pm$ 1.29 87.43 $\pm$ 1.23 86.13
DaDiL-E montesuma2023multi - 90.45 $\pm$ 1.02 77.08 $\pm$ 1.21 86.79 $\pm$ 2.14 89.01 $\pm$ 1.35 84.04 $\pm$ 3.16 87.85 $\pm$ 1.06 85.87
DaDiL-R montesuma2023multi - 91.97 $\pm$ 1.22 77.15 $\pm$ 1.32 85.41 $\pm$ 1.69 89.39 $\pm$ 1.03 84.49 $\pm$ 1.95 88.44 $\pm$ 1.29 86.14
GMM-DaDiL montesuma2024lighter - 91.72 $\pm$ 1.41 76.41 $\pm$ 1.89 89.68 $\pm$ 1.49 89.18 $\pm$ 1.17 86.05 $\pm$ 1.46 88.02 $\pm$ 1.12 86.85
Discrete cuturi2014fast 90.38 $\pm$ 1.29 71.17 $\pm$ 1.26 85.27 $\pm$ 1.09 87.29 $\pm$ 0.99 83.83 $\pm$ 1.74 84.94 $\pm$ 1.87 83.81
NormFlow visentin2025computing 87.06 $\pm$ 0.57 66.25 $\pm$ 1.11 88.58 $\pm$ 1.00 89.35 $\pm$ 0.34 80.99 $\pm$ 2.21 85.08 $\pm$ 1.45 82.89
CW2B korotin2021continuous 91.72 $\pm$ 0.79 73.71 $\pm$ 1.42 88.37 $\pm$ 1.31 89.32 $\pm$ 1.14 84.66 $\pm$ 1.75 87.19 $\pm$ 1.24 85.83
NOT kolesov2024estimating 90.31 $\pm$ 1.08 72.37 $\pm$ 0.94 87.99 $\pm$ 1.31 89.28 $\pm$ 0.81 84.35 $\pm$ 1.68 85.29 $\pm$ 2.02 84.93
U-NOT gazdieva2024robust 90.82 $\pm$ 0.65 72.68 $\pm$ 1.58 88.65 $\pm$ 0.95 89.00 $\pm$ 1.28 85.22 $\pm$ 1.07 86.19 $\pm$ 1.70 85.43
WGF (ours) 91.51 $\pm$ 1.32 73.77 $\pm$ 1.24 87.44 $\pm$ 1.16 89.38 $\pm$ 0.62 84.87 $\pm$ 1.75 86.08 $\pm$ 2.23 85.51
Discrete montesuma2023multi 92.38 $\pm$ 0.66 73.74 $\pm$ 1.07 88.89 $\pm$ 0.85 89.38 $\pm$ 1.26 85.53 $\pm$ 1.35 86.60 $\pm$ 1.63 86.09
GMM montesuma2024lighter 92.23 $\pm$ 0.70 71.81 $\pm$ 1.78 84.72 $\pm$ 1.92 89.28 $\pm$ 1.55 87.51 $\pm$ 1.73 82.49 $\pm$ 1.81 84.67
WGF (ours) 92.34 $\pm$ 1.05 76.10 $\pm$ 1.56 89.78 $\pm$ 0.90 89.32 $\pm$ 1.03 85.92 $\pm$ 1.35 87.74 $\pm$ 1.52 86.87
Table 5: Classification accuracy per domain on the Tennessee Eastman Process benchmark.
We report our results in Table 5. For this benchmark, the gap between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) barycenter methods is closer. For instance, CW2B korotin2021continuous achieves 85.83 for the average target domain performance versus 86.09 for the labeled discrete barycenter. Furthermore, it surpasses the GMM method of montesuma2024lighter. In comparison, our methods establish a new state-of-the-art for barycenter-based domain adaptation in this benchmark, having similar performance to GMM-DaDiL montesuma2024lighter.
Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 Avg. $\uparrow$
CBraMod - 52.22 40.45 64.23 40.41 47.74 48.34 53.64 50.52 55.03 50.30
DaDiL-R montesuma2023multi - 58.68 40.27 66.14 41.45 59.54 51.38 53.12 57.46 55.21 53.69
GMM-DaDiL montesuma2024lighter - 60.59 42.36 70.31 48.12 61.11 50.34 62.15 58.85 60.00 57.10
Discrete cuturi2014fast 60.93 43.40 69.96 50.00 61.28 50.69 61.63 61.45 57.11 57.38
NormFlow visentin2025computing 57.11 43.22 67.53 47.71 60.24 51.04 56.25 60.76 55.20 55.45
CW2B korotin2021continuous 60.00 43.57 70.13 48.95 60.93 51.21 61.28 61.63 57.46 57.25
NOT kolesov2024estimating 61.28 43.22 70.31 47.91 60.59 51.21 61.46 61.80 57.29 57.23
U-NOT gazdieva2024robust 60.59 44.09 69.44 47.70 61.11 50.34 61.97 61.11 58.15 57.17
WGF (ours) 60.93 43.75 69.96 49.16 60.76 51.38 61.97 63.54 57.64 57.68
Discrete montesuma2023multi 60.76 42.53 70.66 48.54 61.28 50.69 61.63 61.28 59.54 57.43
GMM montesuma2024lighter 60.59 42.53 69.79 47.08 61.45 50.34 62.32 59.54 59.72 57.04
WGF (ours) 60.93 46.18 68.92 50.21 62.32 50.69 61.63 61.11 57.11 57.68
Table 6: Classification accuracy per domain on the BCI-CIV-2a benchmark.
BCI-CIV-2a brunner2008bci Contains electroencephalogram (EEG) data from nine subjects. In brief, an EEG measures the electrical activity of the brain over time. As such, like the TEP benchmark, this benchmark has a collection of time series data. The BCI-CIV-2a data is concerned with motor imagery tasks, that is, the imagination of movement of the left hand, right hand, both feet, and tongue. These correspond to class 1, 2, 3, and 4. Therefore, in this benchmark we have four classes. Our goal here is to perform cross-subject adaptation, namely, we use data from a given set of source subjects, and try to predict on a target subject. Therefore, each domain is a subject, so there are nine domains in total. There are, on average, $560$ samples per subject, with a total of $5088$ samples.
We show our results in Table 6. For cross-subject studies, we highlight that, besides improving individual domain performance, it is desirable to not degrade over the source-only performance. This is indeed the case for almost all methods. For instance, DaDiL montesuma2023multi degrades performance of subjects 2 and 7 by a small margin. For this benchmark, the difference between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) is not as marked as before. We hypothesize that this benchmark falls into the covariate shift hypothesis, namely, $Q_{S}(X)\neq Q_{T}(X)$, rather than the more general joint shift $Q_{S}(X,Y)\neq Q_{T}(X,Y)$.
Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
CBraMod - 78.97 80.35 70.46 85.59 80.69 72.85 77.00 73.29 83.41 68.95 76.63
WJDOT turrisi2022multi - 80.41 78.25 65.69 82.85 81.74 77.74 76.33 68.75 86.95 70.81 76.95
DaDiL-E montesuma2023multi - 76.12 77.90 65.93 83.21 81.28 68.45 75.66 72.61 79.75 65.93 75.14
DaDiL-R montesuma2023multi - 78.06 78.13 66.97 84.64 81.51 72.02 76.33 72.84 80.85 67.55 75.89
GMM-DaDiL montesuma2024lighter - 76.63 78.95 66.51 82.73 77.44 73.81 77.00 71.70 83.05 66.86 75.47
Discrete cuturi2014fast 78.77 80.58 69.76 85.83 79.06 73.45 78.33 72.50 84.14 68.48 77.09
NormFlow visentin2025computing 81.12 82.90 75.58 84.28 80.34 78.69 75.88 70.34 80.48 70.81 78.05
CW2B korotin2021continuous 79.08 78.25 69.07 86.54 78.48 70.35 76.88 70.90 82.80 66.04 75.84
NOT kolesov2024estimating 78.57 79.88 67.32 85.59 78.72 73.21 77.77 72.50 83.41 67.21 76.44
U-NOT gazdieva2024robust 80.20 83.48 73.72 87.50 76.51 72.62 76.66 70.22 78.17 71.51 77.06
WGF (ours) 79.69 81.86 70.58 86.66 78.14 77.74 78.77 72.16 85.97 70.00 78.16
Discrete montesuma2023multi 78.87 80.69 72.09 86.07 78.60 78.45 78.00 72.95 85.48 70.81 78.20
GMM montesuma2024lighter 75.20 75.46 69.88 80.71 78.60 71.90 76.33 68.41 78.17 68.14 74.28
WGF (ours) 80.10 80.81 76.62 85.95 76.86 80.35 76.55 77.72 84.87 73.95 80.02
Table 7: Classification accuracy per domain on the ISRUC benchmark.
ISRUC khalighi2013automatic is a benchmark for EEG data for sleep staging. In brief, sleep staging is the classification of physiological data, such as EEG, recorded during sleep into different sleep stages. In this context, the ISRUC benchmark is the most large scale benchmark used in this paper. This study comprises 100 subjects, with a total of 89240 samples (approximately 900 samples per domain). The sleep staging problem in this benchmark has five classes: awake, non-rapid eye movement sleep (subdivided into N1, N2, N3), and rapid eye movement sleep.
Due the elevated number of subjects in this benchmark, we adopt a slightly different evaluation protocol. Instead of using the leave-of-domain-out strategy, we fix the first $90$ subjects as training subjects and evaluate the performance of adapting to the remaining $10$ subjects. This means that we compute Wasserstein barycenters of $K=90$ input measures. Overall, our WGF method has clear advantage over all other methods, with unsupervised WGF (78.16% average target accuracy) being the best over unlabeled barycenters, and labeled barycenters as well (80.02% average test accuracy).
F.3 Complexity Analysis and Additional Running Time
We start with analyzing the complexity of evaluating the functionals used in our work. We recall that,
where, assuming that the complexity of evaluating $V$ and $U$ is $\mathcal{O}(1)$, the complexity of evaluating $\mathbb{V}$ is $\mathcal{O}(n)$ and for $\mathbb{U}$ is $\mathcal{O}(n^{2})$. The complexity of $\mathbb{B}$ is usually the bottleneck, as it involves computing $K$ OT problems. For exact OT, this means we need to solve $K$ OT problems of size $m\times n$, where $m$ is the batch size and $n$ is the barycenter support size. These OT problems need to be solved sequentially, as parallelizing $K$ independent linear programs is non-trivial. In practice, we use the network simplex solver of bonneel2011displacement, wrapped in Python by flamary2021pot. This algorithm has worst-case complexity of $\mathcal{O}(n^{3})$ but in practice, runs in $\mathcal{O}(n^{2})$. Assuming $n\gg m$, the complexity becomes,
$\text{Exact barycenter functional complexity}=\mathcal{O}(K\times m\times n),$
which scales linearly with the barycenter support size $n$. Though this might appear contradictory with the usual $\mathcal{O}(n^{3})$ complexity of OT, we highlight that we are keeping $m$ constant as we scale the barycenter support size $n$. The same analysis can be done for entropic OT,
$\text{Entropic barycenter functional complexity}=\mathcal{O}(K\times L\times m\times n),$
where, instead of a factor $m^{2}$, we have $L\times m$, where $L>0$ is the number of Sinkhorn iterations. Both of these approaches scale linearly with the number of samples in the barycenter support. However, there are 2 important advantages of using entropic OT. First, we can use GPU acceleration for computing the OT variables. Second, as we shown in Section 3.1, we can vectorize the $K$ OT problems. While this does not alleviate the $K$ term in the complexity, it does lead to further speedups. These factors explain the difference, in running time, between exact and entropic OT.
(a) Scaling with $n$ ($\epsilon=0$)
(b) Scaling with $n$ ($\epsilon>0$)
(c) Scaling with $\epsilon$ ($m=256$)
Figure 8: Running time analysis in the Swiss roll example. In (a) and (b), we show that our WGF approach improves the scaling of the discrete solver of cuturi2014fast. In (c), we isolate the speedups of GPU parallelism and Sinkhorn vectorization.
Next, we perform a running time analysis of the tested methods on the ISRUC benchmark, which contains the most number of samples and domains. For all neural net methods, we fix a batch size of $256$ samples. Overall, our proposed methods run significantly faster than previous methods. For instance, in comparison with the discrete solver of cuturi2014fast, our algorithm is approximately seven times faster per iteration. It is noteworthy that some neural net methods are relatively faster per iteration (e.g. NormFlow). However, we empirically verified that these methods sometimes need a large number of iterations to converge to a good barycenter (e.g. 5000 iterations).
Discrete cuturi2014fast, montesuma2023multi $Kn^{3}\log n$ $n$ 4.28 50 214.37
NormFlow visentin2025computing N/A $|f|$ 0.17 5000 878.32
CW2B korotin2021continuous N/A $K(|f|+|g|)$ 2.18 1000 2181.00
NOT kolesov2024estimating N/A $K(|f|+|g|)$ 41.63 25 1040.75
U-NOT gazdieva2024robust N/A $K(|f|+|g|)$ 13.29 100 1329.77
WGF (ours, $\epsilon=0$) $Km^{2}n$ $n$ 0.05 200 10.43
WGF (ours, $\epsilon=10^{-3}$) KLmnKLmn $n$ 0.03 200 7.00
Table 8: Complexity and running time (in seconds) of different barycenter calculation strategies on the ISRUC benchmark. $n_{b}$ denotes the batch size, $n$ number of samples, $K$ number of input measures, $d$ number of dimensions. For NormFlow, $f$ denotes the neural net implementing the normalizing flow. For CW2B, NOT and U-NOT, $f$ and $g$ denote the neural networks approximating the Kantorovich potentials. $|f|$ and $|g|$ denote the number of their parameters.
F.4 Effect of Different Functionals
(a) $(10^{-3},10^{-3})$
(b) $(10^{-4},10^{-3})$
(c) $(10^{-5},10^{-3})$
(d) $(10^{-3},0.0)$
(e) $(10^{-4},0.0)$
(f) $(10^{-5},0.0)$
Figure 9: Wasserstein barycenter as a function of $(\eta_{G},\epsilon)$, i.e., diffusion coefficient $\xi$, and entropy regularization $\epsilon$.
On the effect of $\mathbb{G}$. The internal energy functional [chizat2025doubly] controls the smoothness of $\mathbb{G}$, and acts, in the case of the entropy functional $G(s)=s\log s$, as a diffusion term. In Figure 9, we show the effect of varying $\xi$ (c.f. equation 16) in the final barycenter. Overall, an aggressive diffusion coefficient leads to fuzzy barycenters (c.f. Figure 9 (a)), while a coefficient too small leads to low variance measures (c.f. Figure 9 (c)). This is, in fact, an artifact of the entropic OT. While a large diffusion coefficient does lead to fuzzy barycenters with exact OT (c.f. Figure 9 (d)), using a small coefficient does not degrade the quality of the computed barycenter.
(a) Unsupervised
(b) Unsupervised
(c) Supervised
(d) Supervised
Figure 10: t-SNE embeddings of unsupervised (a, b) and supervised (c, d) barycenters. While it is possible to group some samples according their pseudo-labels based on $\mathbb{U}$, the task is in general difficult. For the supervised case, the interaction energy is able to make classes more separable (d, in comparison to c).
On the effect of $\mathbb{U}$. We now visualize the effect of adding the repulsion interaction energy functional in the Office-Home benchmark. For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 22 and we fix $\text{margin}=1$, which encourages class clusters to live in orthogonal sub-spaces of the latent space. Overall, adding $\mathbb{U}$ leads to well separated classes.
Figure 11: Distribution of label entropy values as a function of potential energy scaling $\eta$.
In Figure 10, we show a comparison between unsupervised and supervised regularized barycenters. In (a), unsupervised, unregularized barycenters do not have any label structure, since there is no label supervision. In (b), we show that by adding the interaction energy functional, we can cluster samples with the same assigned label – even without input measure label supervision. In (c) and (d), we show the supervised case, in which labels in the input measures are available. In (c), classes are well clustered, but some of them overlap in the embedding space. Adding the interaction energy functional $\mathbb{U}$ effectively pushes these apart, as shown in (d). Overall, we conclude that enforcing well separated class clusters is not straightforward without the inductive bias coming from the labels.
On the effect of $\mathbb{V}$. In Figure 11, we show the distribution of label entropy values (c.f. equation 22) as a function of the potential scaling $\eta$,
$\mathbb{V}(P)=\frac{\eta}{n}\sum_{i=1}^{n}V(y_{i}^{(P)}).$
We see that increasing the weight of this functional leads to increasingly sharp labels.
F.5 Hyperparameter Setting
  • number of samples in the barycenter support, $n$.
  • learning rate $\alpha$ for the gradient flow,
  • number of iterations $n_{\text{iter}}$,
[…]